Alarmed at the rising levels of mercury in your town, you decide to estimate the average level in the water supply. From your sample of 25 locations, you calculate a 90% confidence interval of (1.99, 2.11). What is the margin of error in estimating the population mean?

(A) 2.05 (B) 0.06 (C) 1.06 (D) 0.12 (7)

Alarmed at the rising levels of mercury in your town, you decide to estimate the average level in the water supply. You take samples from 25 locations in town. How many degrees of freedom (df) should you use in subsequent hypothesis testing? (A) 5 (B) 24 (C) 25 (D) 26
Answer #1

Solution :

a )Given that,

Lower confidence interval = 2.11

Upper confidence interval = 1.99

\bar x  = (Lower confidence interval + Upper confidence interval ) / 2

= (2.11 + 1.99 ) / 2

= 4.1 / 2

= 2.05

\bar x = 2.05

Margin of error =  E = Upper confidence interval - \bar x

= 2.11 - 2.05

= 0.06

Margin of error =  E = 0.06

Option B ) is correct.:

b )Given that,

n = 25

Degrees of freedom = df = n - 1

= 25 - 1

= 24

Degrees of freedom =24

Option B ) is correct.: